2. A recessive allele (e) is epistatic over the B gene locus, producing a golden cot homozygous (ee). Dogs that are homozygous dominant (EE) o determined by the B locus, either black (BB or Bb) or chocolate (bb). Answer the following qs regarding a cross a Labrador retriever with the genotype Bb Ee and a Labrador retriever with the genotype Bb ee r heterozygous (Ee) will show the color A) What is the phenotype of dog 1? B) What is the phenotype of dog 2? C) What percent of the puppies will be: 1. Black? 2. Brown? 3. Golden?
Answer #1

A) Phenotype of dog 1(Bb Ee) is black,as it does not have ee,so its phenotype is determined by B locus.

B) Phenotype of dog 2(Bb ee) is golden,as it contains ee(e is epistatic over B)

C) After crossing,out of 16 individuals,8 contain ee genotype,so all of them are golden,so golden phenotypes are=(8*100)/16=50%

6 individuals contain either BB or Bb(and do not contain ee).So they are of black phenotypes=(6*100)/16=37.5%

2 individuals contain bb(and do not contain ee).So they are of choclate or brown phenotypes=(2*100)/16=12.5%